# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v2

@@ Abstract
InvestESG is a novel multi-agent reinforcement learning (MARL) benchmark designed to study the impact of Environmental, Social, and Governance (ESG) disclosure mandates on corporate climate investments [^1] Github repo: https://github.com/yuanjiayiy/InvestESG. Supported by both PyTorch and JAX implementation, the benchmark models an intertemporal social dilemma where companies balance short-term profit losses from climate mitigation efforts and long-term benefits from reducing climate risk, while ESG-conscious investors attempt to influence corporate behavior through their investment decisions, in a scalable and hardware-accelerated manner. Companies allocate capital across mitigation, greenwashing, and resilience, with varying strategies influencing climate outcomes and investor preferences. Our experiments show that without ESG-conscious investors with sufficient capital, corporate mitigation efforts remain limited under the disclosure mandate. However, when a critical mass of investors prioritizes ESG, corporate cooperation increases, which in turn reduces climate risks and enhances long-term financial stability. Additionally, providing more information about global climate risks encourages companies to invest more in mitigation, even without investor involvement. Our findings align with empirical research using real-world data, highlighting MARL’s potential to inform policy by providing insights into large-scale socio-economic challenges through efficient testing of alternative policy and market designs.
@@ 1 Introduction
Climate change poses a persistent threat to global stability, with droughts, storms, fires, and flooding becoming more intense and frequent (Christopher B Field and Dahe, 2012), leading to disruption of the natural ecosystem and significant impacts on the global economy. Addressing climate change requires coordinated efforts across multiple sectors, particularly from large corporations, which are reportedly responsible for over 70% of global industrial greenhouse gas emissions (Griffin and Heede, 2017). While adaptation—preparing for the inevitable consequences of climate change—tends to be party-specific and often driven by financial incentives, mitigation—reducing emissions—presents a social dilemma (Leibo et al., 2017), where the benefits of reduced emissions are shared globally yet the costs are borne locally (Olson Jr, 1971; Dahlman, 1979; Buchanan and Stubblebine, 2006). As corporations are inherently self-interested, they are unlikely to reduce emissions voluntarily without external incentives or regulations.
Numerous policies have been proposed to address this challenge. Among these, mandatory Environmental, Social, and Governance (ESG) disclosures have recently been hotly debated in the United States. The Securities and Exchange Commission’s (SEC) proposal, which would require publicly traded companies to disclose climate-related risks, mitigation strategies, and greenhouse gas emissions from their operations, has attracted over 15,000 comments, making it one of the most contentious proposals in the SEC’s history (SEC, 2024a; CNBC, 2024). This has resulted in an indefinite delay in enactment of the policy to allow for further discussion (SEC, 2024b). The U.S. is not alone in facing such pushback; similar delays are unfolding in the European Union and Korea (Bloomberg News, 2024; Korea Economic Daily, 2023). This highlights the need for thorough research to effectively inform the design and implementation of these policies.
Traditional economics and policy research relies on either empirical analysis—which does not enable testing possible new policies (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021)—or theoretical economics models, which are often limited to scenarios with only two agents (e.g., (Friedman et al., 2021)), or single-period games (e.g., (Pázstor et al., 2021)). In contrast, Multi-Agent Reinforcement Learning (MARL) enables simulating complex interactions between multiple agents over extended time periods, under diverse hypothesized policy settings. Leveraging MARL to address large-scale socio-economic questions is a growing field (Hertz et al., 2023). Prior work has demonstrated the potential of MARL to design effective taxation schemes that enhance both equality and productivity (Zheng et al., 2021), highlighting its relevance for tackling social challenges.
@@
Our aim is not to claim that our model fully represents real-world complexity. It remains a "first-principles" model; however, it captures aspects that are challenging to integrate into a single framework using traditional economic approaches. We present InvestESG as a challenging benchmark for the machine learning community; for researchers that are interested in developing MARL algorithms that can solve social dilemmas, InvestESG represents a social dilemma environment that could inform policymaking. We will provide the code for the benchmark and agent baselines in open-source.
@@ 2 Related Work
Conventional economic methods are limited by either generalizability or tractability. Existing ESG disclosure related research in economics, business, and public policy rely on either empirical data (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021) or simplified theoretical models (Polinsky and Shavell, 2012; Kalkanci and Plambeck, 2020; Cho et al., 2019; Pázstor et al., 2021; Friedman et al., 2021). Empirical analyses, while grounded in real-world data, struggle with generalizability and testing counterfactual policies. Theoretical models provide formalized equilibria and explore counterfactuals but are limited by tractability, modeling either multiple agents in single-period games (e.g., (Pázstor et al., 2021)) or only two agents over limited time periods (e.g., (Friedman et al., 2021)). By proposing a MARL framework, our method overcomes these limitations by enabling the simulation of complex, multi-agent systems over extended time horizons under diverse policy settings, which allows for emergent behaviors and better captures complex socio-economic dynamics among diverse agents (Hertz et al., 2023).
@@
Current MARL benchmarks and social dilemma environments were not designed to model specific policy problems. Various MARL benchmarks have been created to study multi-agent coordination and cooperation; however, they are often limited to simplistic particle simulations (Lowe et al., 2017) or videogames (Agapiou et al., 2023; Carroll et al., 2020; Samvelyan et al., 2019; Bettini et al., 2024a, b), which have little direct real-world implication. Sequential social dilemmas (SSD) (Leibo et al., 2017) are spatially and temporally extended multi-agent environments in which the payoff to an individual agent for defecting is higher, but if all agents defect the payoff is lower. SSDs go beyond traditional game-theoretic environments like Prisoner’s Dilemma because the complexity of the solving the SSD depends on not only addressing the misalignment between individual and collective rationality, but doing so when the negative consequences of short-sighted actions may take a long time to manifest. Prior research has examined methods to promote cooperation in SSDs by incorporating inequity aversion in agents (Hughes et al., 2018), rewarding agents for influencing others’ actions (Jaques et al., 2018), enabling agents to provide incentives to others (Yang et al., 2020a), and controlling for behavioral diversity (Bettini et al., 2024b). Many of these studies are inspired by factors that drive human cooperation in social dilemmas. However, much of this research has been conducted in environments with no direct real-world implications (e.g. (Leibo et al., 2017, 2021)). In contrast, our environment directly addresses the problem of climate change, and is designed with critical trade-offs around a certain policy-making question. We hope to encourage researchers interested in addressing social dilemmas to focus on an environment with potential impact on the problem of climate change (Bisaro and Hinkel, 2016).
@@ 3 Problem Setting
Corporations face two key trade-offs in climate action. The first is between short-term costs and long-term benefits, as climate efforts often require upfront investments with returns realized over time. The second is between self-interest and collective benefit, as emission reductions benefit all, but the costs are borne by those who act, which incentivizes free-riding and discourages collective action. These trade-offs motivate us to conceptualize corporate climate actions as an intertemporal social dilemma (Leibo et al., 2017; Bisaro and Hinkel, 2016), which by definition, involves a conflict between immediate self-interest and longer-term collective interests.
We introduce InvestESG, a MARL benchmark environment designed to evaluate the impact of ESG disclosure mandates on corporate climate investments. The environment is grounded in well-established finance and economics literature that explores similar questions, providing a conceptual framework for studying corporate and investor dynamics in response to ESG policies. Pastor et al. (2021) and Pedersen et al. (2021) proposed analytical models of single-period equilibria, where companies either choose or are assigned fixed levels of "greenness," and investors optimize portfolios to balance financial returns with preferences for green assets. We follow a similar setting where companies choose mitigation effort levels every period and receive a single metric ESG score that reflects their greenness. Investors derive utility from the profits generated by their portfolios and, based on their level of "ESG-consciousness", may also receive positive utility by investing in firms with high ESG scores. We extend these frameworks by (1) modeling the evolution of firm-investor strategies over the long run, which poses significant analytical and numerical challenges (Pakes and McGuire, 2001), (2) incorporating climate evolution, where risks worsen over time but can be mitigated by firm efforts, and (3) introducing greenwashing (Lyon and Maxwell, 2011) where companies cheaply inflate ESG scores without genuine mitigation, which raises issues of information asymmetry and principal-agent problems (Grossman and Hart, 1992).
We aim to address the following research questions: (1) Do ESG-conscious investors incentivize companies to undertake mitigation efforts, and how does the level of ESG-consciousness influence these outcomes? (2) Does the presence of heterogeneous investors lead to bifurcation in agent strategies, where only some companies engage in mitigation and attract investments from ESG-conscious investors, while others and their investors opt to free-ride? (3) Are companies likely to engage in greenwashing? (4) What measures can enhance the effectiveness of ESG disclosure mandates?
@@ 4 The InvestESG Environment
In this section, we provide a brief overview of the environment setup. The detailed mathematical formulation is presented in Appendix 9.
The simulation starts in 2021 and runs through 2120, with each period $t$ corresponding to one year. The environment includes two main components: (1) an evolving climate and economic system, and (2) two types of agents: $M$ company agents $\mathcal{C}_{i}$ for $i\in\{1,\dots,M\}$ and $N$ investor agents $\mathcal{I}_{j}$ for $j\in\{1,\dots,N\}$.
@@
(b) Average number of severe climate event occurrences over 10-year period.
Figure 2: Status quo scenario where all agents are only profit-motivated. In (a), mitigation spending (blue curve) is minimal, leading climate risk (yellow curve) to increase over time. Adverse weather event occurrences are shown as dotted lines; red lines indicate multiple adverse events in a single year. In (b), increasing climate risk leads to more frequent occurrence of severe climate events over time.
@@
Company Action Space. Each company agent $\mathcal{C}_{i}$ in period $t$ selects actions from a continuous vector $\mathbf{u}_{t}^{\mathcal{C}_{i}}=(u_{t,m}^{\mathcal{C}_{i}},u_{t,g}^{\mathcal{C}_{i}},u_{t,r}^{\mathcal{C}_{i}})$, where $u_{t,m}^{\mathcal{C}_{i}}$ represents the share of capital allocated to mitigation, $u_{t,g}^{\mathcal{C}_{i}}$ to greenwashing, and $u_{t,r}^{\mathcal{C}_{i}}$ to building climate resilience. The action space for each company is defined as a continuous 3-dimensional unit cube, $\mathcal{U}_{t}^{\mathcal{C}_{i}}=[0,1]^{3}$. If the sum of the three ratios exceeds 1 in any period, the company agent is deemed to be overspending, resulting in bankruptcy. Mitigation spending directly reduces the system-wide climate risk as explained above. Greenwashing involves deceptive marketing or accounting tactics that allow companies to appear climate-friendly at a low cost without providing any real benefits to society (de Freitas Netto et al., 2020; Yang et al., 2020b). Resilience spending enhances the company’s climate resilience by lowering its vulnerability $L_{t}^{\mathcal{C}_{i}}$, but it does not reduce emissions and therefore does not mitigate system-wide climate risk. In the default setting, we disable greenwashing and resilience spending to focus on testing companies’ mitigation efforts. These options are later enabled to examine their effects, as will become clear in Section 5.
@@
Investor Action Space. Investor agents first select the set of companies they want to invest in. Specifically, agent $\mathcal{I}_{j}$ in period $t$ selects an action from a binary vector of length $M$, $\mathbf{a}_{t}^{\mathcal{I}_{j}}=(a_{t,1}^{\mathcal{I}_{j}},\dots,a_{t,M}^{\mathcal{I}_{j}})$, corresponding to the $M$ company agents. Each entry $a_{t,i}^{\mathcal{I}_{j}}=1$ indicates that investor $\mathcal{I}_{j}$ invests in company $\mathcal{C}_{i}$ in period $t$, and $a_{t,i}^{\mathcal{I}_{j}}=0$ otherwise. The investor’s action space at time $t$ is thus $\mathcal{A}_{t}^{\mathcal{I}_{j}}=\{0,1\}^{M}$. Once the choices are made, investors capitals are distributed equally among these chosen companies.
@@
Modeling ESG Disclosure. With the ESG disclosure mandate in place, each company agent receives an updated ESG score $Q_{t+1}^{\mathcal{C}_{i}}$ in period $t$, calculated as $Q_{t+1}^{\mathcal{C}_{i}}=u_{t,m}^{\mathcal{C}_{i}}+\beta u_{t,g}^{\mathcal{C}_{i}}$, where $\beta>1$ indicates that greenwashing is cheaper than genuine mitigation in terms of building an ESG-friendly image. This mirrors simple ESG ratings provided by some agencies, which typically range from 1 to 100 (S&P Global, 2024) or use letter-based ratings (MSCI Inc., 2024).
@@
State and Observation Space. The environment simulates a partially observable Markov game $\mathcal{M}$ defined over a continuous, multi-dimensional state space. The system state at period $t$ is characterized by the climate risk parameter $P_{t}$, each company agent’s state vector $\mathbf{S}_{t}^{\mathcal{C}_{i}}=(K_{t}^{\mathcal{C}_{i}},Q_{t}^{\mathcal{C}_{i}},L_{t}^{\mathcal{C}_{i}})$, where $K_{t}^{\mathcal{C}_{i}}$ is the capital level, $Q_{t}^{\mathcal{C}_{i}}$ is the ESG score, and $L_{t}^{\mathcal{C}_{i}}$ is the climate resilience. Each investor agent’s state is represented by their investment portfolio and cash levels $\mathbf{S}_{t}^{\mathcal{I}_{j}}=(H_{t,1}^{\mathcal{I}_{j}},\dots,H_{t,M}^{\mathcal{I}_{j}},C_{t}^{\mathcal{I}_{j}})$, where $H_{t,i}^{\mathcal{I}_{j}}$ represents investor $\mathcal{I}_{j}$’s holdings in company $\mathcal{C}_{i}$, and $C_{t}^{\mathcal{I}_{j}}$ is the investor’s cash level. The full system state at period $t$ is thus $\mathcal{S}_{t}=\left(P_{t},\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. All company and investor agents share a common observation space, denoted as $\mathcal{O}_{t}$. In the default setting, $\mathcal{O}_{t}=\left(\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. Extensions that incorporate additional observable information, like climate risk, are explored in Section 5.
@@
State Transition. The environment’s state transition $\mathcal{T}$ proceeds as follows. At the beginning of period $t$, investors collect their investment holdings from period $t-1$ and redistribute their capital according to $\mathbf{a}_{t}^{\mathcal{I}_{j}}$. If an investor chooses not to invest, all capital remains as cash. Companies then make climate-related spending, after which system climate risk and companies’ inidividual resilience are updated. Meanwhile, companies receive updated ESG scores, and the occurrence of climate events are simulated. Afterwards, companies’ profit margins, $\rho_{t}^{\mathcal{C}_{i}}$, are computed as a function of their climate-related spending $\mathbf{u}_{t}^{\mathcal{C}_{i}}$, default economic growth $\gamma$, and losses due to climate events, $L_{t}^{\mathcal{C}_{i}}$, and the number of climate events $X_{t}$. Finally, company capitals are updated, so are investor holdings and cash positions based on the profit margins of their portfolios.

### v3

@@ Abstract
InvestESG is a novel multi-agent reinforcement learning (MARL) benchmark designed to study the impact of Environmental, Social, and Governance (ESG) disclosure mandates on corporate climate investments. The benchmark models an intertemporal social dilemma where companies balance short-term profit losses from climate mitigation efforts and long-term benefits from reducing climate risk, while ESG-conscious investors attempt to influence corporate behavior through their investment decisions. Companies allocate capital across mitigation, greenwashing, and resilience, with varying strategies influencing climate outcomes and investor preferences. We are releasing open-source versions of InvestESG in both PyTorch and JAX [^1] Github repo: https://github.com/yuanjiayiy/InvestESG, which enable scalable and hardware-accelerated simulations for investigating competing incentives in mitigate climate change. Our experiments show that without ESG-conscious investors with sufficient capital, corporate mitigation efforts remain limited under the disclosure mandate. However, when a critical mass of investors prioritizes ESG, corporate cooperation increases, which in turn reduces climate risks and enhances long-term financial stability. Additionally, providing more information about global climate risks encourages companies to invest more in mitigation, even without investor involvement. Our findings align with empirical research using real-world data, highlighting MARL’s potential to inform policy by providing insights into large-scale socio-economic challenges through efficient testing of alternative policy and market designs.
@@ 1 Introduction
Climate change poses a persistent threat to global stability, with droughts, storms, fires, and flooding becoming more intense and frequent (Field et al., 2012), leading to disruption of the natural ecosystem and significant impacts on the global economy. Addressing climate change requires coordinated efforts across multiple sectors, particularly from large corporations, which are reportedly responsible for over 70% of global industrial greenhouse gas emissions (Griffin & Heede, 2017). While adaptation—preparing for the inevitable consequences of climate change—tends to be party-specific and often driven by financial incentives, mitigation—reducing emissions—presents a an intertemporal social dilemma (Leibo et al., 2017; Hughes et al., 2018), where the benefits of reduced emissions are shared globally yet the costs are borne locally (Olson Jr, 1971; Dahlman, 1979; Buchanan & Stubblebine, 2006). As corporations are inherently self-interested, they are unlikely to reduce emissions voluntarily without external incentives or regulations.
Numerous policies have been proposed to address this challenge. Among these, mandatory Environmental, Social, and Governance (ESG) disclosures have recently been hotly debated in the United States. The Securities and Exchange Commission’s (SEC) ESG proposal, which would require publicly traded companies to disclose climate-related risks, mitigation strategies, and greenhouse gas emissions from their operations, has attracted over 15,000 comments, making it one of the most contentious proposals in the SEC’s history (SEC, 2024b; CNBC, 2024). This has resulted in an indefinite delay in enactment of the policy to allow for further discussion (SEC, 2024a). The U.S. is not alone in facing such pushback; similar delays are unfolding in the European Union and Korea (Bloomberg News, 2024; Korea Economic Daily, 2023). This highlights the need for thorough research to effectively inform the design and implementation of these policies.
Traditional economics and policy research relies on either empirical analysis—which does not enable testing possible new policies (Doshi et al., 2013; Li & Wu, 2020; Krueger et al., 2021)—or theoretical economics models, which are often limited to scenarios with only two agents (e.g., Friedman et al. 2021), or single-period games (e.g., Pástor et al. 2021). In contrast, Multi-Agent Reinforcement Learning (MARL) enables simulating complex interactions between multiple agents over extended time periods, under diverse hypothesized policy settings. Leveraging MARL to address large-scale socio-economic questions is a growing field (Hertz et al., 2023). Prior work has demonstrated the potential of MARL to design effective taxation schemes that enhance both equality and productivity (Zheng et al., 2021), highlighting its relevance for tackling social challenges.
@@
Our aim is not to claim that our model fully represents real-world complexity. It remains a “first-principles” model; however, it captures aspects that are challenging to integrate into a single framework using a traditional economic approach. We present InvestESG as a challenge for the machine learning community; for researchers interested in developing better cooperative agents, or improved MARL algorithms, InvestESG represents a benchmark that could actually inform policymaking.
@@ 2 Related Work
Conventional economic methods are limited by either generalizability or tractability. Existing ESG disclosure research in economics, business, and public policy rely on either empirical data (Doshi et al., 2013; Li & Wu, 2020; Krueger et al., 2021) or simplified theoretical models (Polinsky & Shavell, 2012; Kalkanci & Plambeck, 2020; Cho et al., 2019; Pástor et al., 2021; Friedman et al., 2021). Empirical analyses, while grounded in real-world data, struggle with generalizability and testing counterfactual policies. Theoretical models provide formalized equilibria and explore counterfactuals but are limited by tractability, modeling either multiple agents in single-period games (e.g., Pástor et al. 2021) or only two agents over limited time periods (e.g., Friedman et al. 2021). By proposing a MARL framework, our method overcomes these limitations by enabling the simulation of complex, multi-agent systems over extended time horizons under diverse policy settings, which allows for emergent behaviors and better captures complex socio-economic dynamics among diverse agents (Hertz et al., 2023).
@@
Current MARL benchmarks and social dilemma environments were not designed to model specific policy problems. Various MARL benchmarks have been created to study multi-agent coordination and cooperation; however, they are often limited to simplistic particle simulations (Lowe et al., 2017) or videogames (Agapiou et al., 2023; Carroll et al., 2020; Samvelyan et al., 2019; Bettini et al., 2024b; a), which have little direct real-world implication. Sequential social dilemmas (SSD) (Leibo et al., 2017) are spatially and temporally extended multi-agent environments in which the payoff to an individual agent for defecting is higher, but if all agents defect the payoff is lower. SSDs go beyond traditional game-theoretic environments like Prisoner’s Dilemma because the complexity of the solving the SSD depends on not only addressing the misalignment between individual and collective rationality, but doing so when the negative consequences of short-sighted actions may take a long time to manifest. Prior research has examined methods to promote cooperation in SSDs by incorporating inequity aversion in agents (Hughes et al., 2018), rewarding agents for influencing others’ actions (Jaques et al., 2018), enabling agents to provide incentives to others (Yang et al., 2020a), and controlling for behavioral diversity (Bettini et al., 2024a). Many of these studies are inspired by factors that drive human cooperation in social dilemmas. However, much of this research has been conducted in environments with no direct real-world implications (e.g. Leibo et al. 2017; 2021). In contrast, our environment directly addresses the problem of climate change, and is designed with critical trade-offs around a specific policy question. We hope to encourage researchers interested in addressing social dilemmas to focus on an environment with potential beneficial social impact on the problem of climate change (Bisaro & Hinkel, 2016).
@@ 3 Problem Setting
Corporations face two key trade-offs in climate action. The first is between short-term costs and long-term benefits, as climate efforts often require upfront investments with returns realized over time. The second is between self-interest and collective benefit, as emission reductions benefit all, but the costs are borne by those who act, which incentivizes free-riding and discourages collective action. These trade-offs motivate us to conceptualize corporate climate actions as an intertemporal social dilemma (Leibo et al., 2017; Bisaro & Hinkel, 2016), which by definition, involves a conflict between immediate self-interest and longer-term collective interests.
We introduce InvestESG, a MARL benchmark environment designed to evaluate the impact of ESG disclosure mandates on corporate climate investments. The environment is grounded in well-established finance and economics literature that explores similar questions, providing a conceptual framework for studying corporate and investor dynamics in response to ESG policies. Pastor et al. (2021) and Pedersen et al. (2021) proposed analytical models of single-period equilibria, where companies either choose or are assigned fixed levels of “greenness,” and investors optimize portfolios to balance financial returns with preferences for green assets. We follow a similar setting where companies choose mitigation effort levels every period and receive a single metric ESG score that reflects their greenness. Investors derive utility from the profits generated by their portfolios and, based on their level of “ESG-consciousness”, may also receive positive utility by investing in firms with high ESG scores. We extend these frameworks by (1) modeling the evolution of firm-investor strategies over the long run, which poses significant analytical and numerical challenges (Pakes & McGuire, 2001), (2) incorporating climate evolution, where risks worsen over time but can be mitigated by firm efforts, and (3) introducing greenwashing (Lyon & Maxwell, 2011) where companies cheaply inflate ESG scores without genuine mitigation, which raises issues of information asymmetry and principal-agent problems (Grossman & Hart, 1992).
We aim to address the following research questions: (1) Do ESG-conscious investors incentivize companies to undertake mitigation efforts, and how does the level of ESG-consciousness influence these outcomes? (2) Does the presence of heterogeneous investors lead to bifurcation in agent strategies, where only some companies engage in mitigation to attract ESG-conscious investors, while others and their investors opt to free-ride? (3) Are companies likely to engage in greenwashing? (4) What measures can enhance the effectiveness of ESG disclosure mandates?
@@ 4 The InvestESG Environment
In this section, we provide a brief overview of the environment. The detailed mathematical formulation is presented in Appendix 9. The simulation starts in 2021 and runs through 2120, with each period $t$ corresponding to one year. The environment includes two main components: (1) an evolving climate and economic system, and (2) two types of agents: $M$ company agents $\mathcal{C}_{i}$ for $i\in\{1,\dots,M\}$ and $N$ investor agents $\mathcal{I}_{j}$ for $j\in\{1,\dots,N\}$.
@@
(b) Average number of severe climate event occurrences over 10-year period for the scenario in part (a).
Figure 2: Status quo scenario where all agents are only profit-motivated. In (a), mitigation spending (blue curve) is minimal, leading climate risk (yellow curve) to increase over time. Adverse weather event occurrences are shown as dotted lines; red lines indicate multiple adverse events in a single year. (b) plots the average number severe climate events over the episode in (a), showing how increasing climate risk leads to more frequent extreme weather events.
@@
Action Space. Each company agent $\mathcal{C}_{i}$ in period $t$ selects actions from a continuous vector $\mathbf{u}_{t}^{\mathcal{C}_{i}}=(u_{t,m}^{\mathcal{C}_{i}},u_{t,g}^{\mathcal{C}_{i}},u_{t,r}^{\mathcal{C}_{i}})$, where $u_{t,m}^{\mathcal{C}_{i}}$ represents the share of capital allocated to mitigation, $u_{t,g}^{\mathcal{C}_{i}}$ to greenwashing, and $u_{t,r}^{\mathcal{C}_{i}}$ to building climate resilience.
@@
Each investor agent first selects the set of companies it will invest in. Specifically, agent $\mathcal{I}_{j}$ in period $t$ selects an action from a binary vector of length $M$, $\mathbf{a}_{t}^{\mathcal{I}_{j}}=(a_{t,1}^{\mathcal{I}_{j}},\dots,a_{t,M}^{\mathcal{I}_{j}})$, corresponding to the $M$ company agents.
@@
Modeling ESG Disclosure. With the ESG disclosure mandate in place, each company agent receives an updated ESG score $Q_{t+1}^{\mathcal{C}_{i}}$ in period $t$, calculated as $Q_{t+1}^{\mathcal{C}_{i}}=u_{t,m}^{\mathcal{C}_{i}}+\beta u_{t,g}^{\mathcal{C}_{i}}$, based on mitigation and greenwash spending, where $\beta>1$ indicates that greenwashing is cheaper than genuine mitigation in terms of building an ESG-friendly image.
@@
State and Observation Space. The environment simulates a partially observable Markov game $\mathcal{M}$ defined over a continuous, multi-dimensional state space. The system state at period $t$ is characterized by the climate risk parameter $P_{t}$, each company agent’s state vector, consistent of its capital level, its ESG, and its climate resilience. Each investor agent’s state is represented by their investment portfolio and cash levels. All company and investor agents share a common observation space, with the state of all companies and investors. Extensions that incorporate additional observable information, like climate risk, are explored in Section 5.
@@
State Transition. At the beginning of period $t$, investors collect their investment holdings from period $t-1$ and redistribute their capital according to $\mathbf{a}_{t}^{\mathcal{I}_{j}}$. If an investor opts not to invest, all capital remains as cash. Companies then make climate-related spending, after which system climate risk and companies’ individual resilience are updated. Meanwhile, companies receive updated ESG scores, and the occurrence of climate events are simulated. Afterwards, companies’ profit margins, $\rho_{t}^{\mathcal{C}_{i}}$, are computed as a function of their climate-related spending $\mathbf{u}_{t}^{\mathcal{C}_{i}}$, default economic growth $\gamma$, and losses due to climate events, $L_{t}^{\mathcal{C}_{i}}$, and the number of climate events $X_{t}$. Finally, company capitals are updated, and so are investor holdings and cash based on the profit margins of their portfolios.